# Chatbot Companionship: A Mixed-Methods Study of Companion Chatbot Usage Patterns and Their Relationship to Loneliness in Active Users

### v2

@@ Abstract
Companion chatbots offer a potential solution to the growing epidemic of loneliness, but their impact on users’ psychosocial well-being remains poorly understood. This study presents a large-scale survey (n = 404) of regular users of companion chatbots, investigating the relationship between chatbot usage and loneliness. We develop a model explaining approximately 50% of variance in loneliness; while usage does not directly predict loneliness, we identify factors including neuroticism, social network size, and problematic use. We identify seven distinct clusters of users, from socially fulfilled dependent users to lonely moderate users. Different usage patterns can lead to markedly different outcomes, with some users experiencing enhanced social confidence while others risk further isolation. Our work contributes to the ongoing dialogue about the role of AI in social and emotional support, offering insights for developing more targeted and ethical approaches to AI companionship that complement rather than replace human connections.
Keywords Companion chatbots, loneliness, artificial intelligence, artificial agents, human-AI interaction, social support
Figure 1: Overview of the study’s methodology and research questions. A survey of 404 participants explored chatbot usage and its psychological impacts. RQ1 investigates the motivations and uses of chatbots through descriptive statistics and free responses. RQ2 examines the relationship between chatbot usage and loneliness using correlation and multiple regression analyses, identifying potential mediating and moderating factors. RQ3 focuses on identifying distinct user profiles through K-means cluster analysis, considering psychological and social factors. The results highlight the complexity of human-AI interactions and their varied impacts on emotional well-being.
@@ 1 Introduction
In an increasingly digital world, we are paradoxically more connected and isolated than ever. Despite unprecedented technological means to connect, loneliness remains pervasive, affecting approximately a third of individuals in industrialized countries [Loveys et al.(2019)]. This phenomenon has led to what the U.S. Surgeon General declared an “epidemic of loneliness”[Office of the Assistant Secretary for Health (OASH)(2023)]. The gravity of this issue is underscored by research indicating that chronic loneliness can increase mortality risk by 26-29%, comparable to smoking 15 cigarettes a day [Chawla et al.(2021)].
Artificial intelligence (AI) has emerged as a potential tool for addressing loneliness. Companion chatbots, which engage in complex conversations and maintain persistent relationships with users, have gained significant popularity, with platforms like Replika and Character.ai reporting over 30 million and 20 million users, respectively [Patel(2024), Shewale(2024)]. These AI companions offer users the opportunity for constant, judgment-free interaction. However, the impact of these artificial relationships on users’ psychological well-being and social behaviors needs further study.
Some research suggests that AI companions, even when used over the short term, decrease feelings of loneliness [Loveys et al.(2019), Gasteiger et al.(2021), De Freitas et al.(2024)]. However, concerns have been raised about the authenticity of these relationships, suggesting that they lower the bar for what people think relationships can be [Turkle(2011)]. This could lead to unrealistic expectations and further social withdrawal.
The nuanced nature of human-AI interaction in this context presents a critical gap in our understanding. Specifically, we lack comprehensive knowledge about factors influencing companion chatbots’ impact on loneliness and other psychological and social factors. As Xia et al. [Xia et al.(2024)] noted, artificial agents may complement human social interactions in some contexts while replacing them in others. This variability in outcomes underscores the need for a more comprehensive and granular analysis of user characteristics, usage patterns, and their relationships to psychological well-being.
To address this gap, we conducted a survey (n = 404) of regular companion chatbot users, with the following research questions:
@@
Our study (see Figure 1) aims to provide a more nuanced understanding of the impact of companion chatbots. By identifying the characteristics of users who experience psychosocial benefits from chatbot interactions (i.e. reduced loneliness and improved social skills) and those who may experience negative consequences (i.e. emotional dependence and social withdrawal), we hope to inform more targeted and ethical design practices. For example, more information could guide the development of customizable interfaces that emphasize features beneficial to specific user groups or implement safeguards for at-risk users. By examining how users’ mental models and expectations of AI companions influence outcomes, we can explore strategies for fostering healthy human-AI relationships that complement rather than replace human connections.
Through a combination of statistical analyses, including multiple regression and cluster analysis, along with qualitative assessment of user responses, we present the following contributions:
  • A model explaining approximately 50% of variance in loneliness, identifying problematic use as a key mediator and social attraction as a moderator in the relationship between chatbot usage and loneliness while accounting for personality traits and social network characteristics.
  • A typology of seven distinct companion chatbot user profiles demonstrating how similar usage patterns can lead to markedly different outcomes based on individual characteristics and social contexts.
  • An exploration of motivations and usage characteristics across different user types, revealing how technological exploration and entertainment dominate over social purposes despite marketing as social tools.
  • Design implications for creating more effective and ethically responsible AI companions, including the need for personalized approaches and built-in mechanisms to detect and intervene in potentially harmful usage patterns.
@@ 2.1 Loneliness and Social Support
Loneliness, defined as the subjective feeling of isolation regardless of one’s actual social network, has become a pressing global issue affecting approximately a third of individuals in industrialized countries [Loveys et al.(2019)]. Social isolation is the corresponding objective measure, often defined as a poorness in quantity and quality of contact with others. The U.S. Surgeon General’s declaration of loneliness as an epidemic in 2023 underscores the gravity of this issue [Office of the Assistant Secretary for Health (OASH)(2023)]. Chronic loneliness is associated with a significantly increased risk of mortality—approximately 26-29%, comparable to smoking 15 cigarettes a day or having an alcohol use disorder [Freedman and Nicolle(2020)].
Various risk factors contribute to social isolation and loneliness, including socio-demographic factors (e.g., greater age, low income), medical factors (e.g., cognitive impairment, sensory loss), and social factors (e.g., living alone, lack of transportation) [Freedman and Nicolle(2020)]. While there is evidence of loneliness peaking in adolescence in certain demographics [Yang and Victor(2011)], older populations are particularly vulnerable due to the accumulation of these risk factors and potentially lower digital literacy, which can limit their use of modern communication technologies [Gasteiger et al.(2021)].
Social support has been identified as a strong predictor of good physical and mental health, as well as resilience to stress. It can be defined as the support accessible to an individual via their social ties to others including individuals, groups, and a larger community [Ozbay et al.(2007)]. Importantly, the quality of relationships is a better predictor of health outcomes than the mere quantity of social connections [Wang et al.(2018), Schutter et al.(2021)].
2.2 Loneliness Coping Strategies and Interventions
Individuals employ various strategies to cope with loneliness, ranging from traditional approaches to digital solutions. Rokach (1998) identified six primary factors in coping with loneliness: Reflection and Acceptance (involving solitary reflection and acceptance of loneliness), Self-Development and Understanding (involving increased understanding of oneself through relationships and therapy), Social Support Network (increased social involvement), Distancing and Denial (involving unhealthy behaviors such as turning to drugs and self-harming), Religion and Faith (involving the feeling of belonging in religious communities), and Increased Activity (devoting more time to extracurricular activities) [Rokach and Brock(1998)]. A systematic review by Deckx et al. (2018) further categorized coping strategies into problem-focused and emotion-focused approaches, with problem-focused strategies consistently associated with lower levels of loneliness [Deckx et al.(2018)].
In recent years, digital platforms have emerged as commonly used coping mechanisms for loneliness, providing both social interaction and passive engagement. Individuals may use video platforms such as YouTube to cope with their loneliness – higher loneliness is associated with the formation of parasocial relationships, where individuals form one-sided relationships with content creators [Niu et al.(2021), de Bérail et al.(2019), Balcombe and Leo(2023)]. This is consistent with the parasocial compensation hypothesis, which postulates that lonely, socially isolated, and socially anxious individuals tend to be more engaged in parasocial relationships [Madison et al.(2016)].
Online communities and forums, such as Reddit, Facebook Groups, and Twitch, facilitate real-time communication and foster social connections across shared interests, allowing individuals to engage in virtual communities and potentially alleviate social isolation [Seering et al.(2019)]. Existing work suggests that seeking and offering social support are key activities in forums [Johnson et al.(2020)]. However, high levels of loneliness and social anxiety are associated with addiction and problematic use of these platforms, mediated by increased social comparison and fear of missing out [Reer et al.(2019), O’Day and Heimberg(2021), Burke et al.(2010)]. Interestingly, image-based media like Instagram may decrease loneliness due to the enhanced intimacy of image-sharing, while text-based media such as X (Twitter) does not show the same effect [Pittman and Reich(2016)].
There are, however, effective interventions for reducing loneliness. These typically fall into four categories, and can be delivered in-person or online: (1) modifying maladaptive social cognition (e.g., cognitive behavioral therapy, psychodynamic therapy, reminiscence, and mindfulness), (2) improving social support (e.g., involving care professionals in health and social care provision), (3) facilitating social interaction (e.g., clubs, shared interest groups, and videoconferencing), and (4) enhancing social skills (e.g., training programs and computer courses) [Rokach and Brock(1998), Gasteiger et al.(2021), Hickin et al.(2021)]. While these interventions can reduce loneliness, their efficacy varies across individuals. Hickin et al. [Hickin et al.(2021)] conducted a systematic review and meta-analysis of psychological interventions for loneliness, finding that such interventions are generally effective across the lifespan. However, they emphasize that the considerable heterogeneity in intervention effectiveness underscores the importance of personalized approaches. This finding aligns with the growing recognition that lonely individuals are a diverse group with varying needs [Cacioppo and Cacioppo(2018)], suggesting that interventions – including technological ones – must be tailored to individual circumstances for optimal impact.
@@ 2.3 Companion Chatbots
The emergence of companion chatbots, also known as virtual friends or digital companions, represents a novel approach to tackling loneliness [Wang et al.(2024), Goodings et al.(2024), Boine(2023), Chen et al.(2024), Shani et al.(2022), Brewer(2022)]. We refer to these as commercially available chatbots that employ generative AI to engage in complex conversations and maintain persistent relationships with users. These companion chatbots are accessible to a broad demographic, with several popular companion chatbot services (see Figure 2) having gained significant user bases:
Figure 2: Examples of companion chatbots: Replika, Character.ai, Snapchat’s My AI, Anima, and Pi.
  • Replika: An app featuring a generative AI chatbot designed to act as a friend, sibling, assistant, or romantic partner. Users interact with one chatbot that adaptively develops over time through continued conversations. Replika’s CEO reported over 30 million registered users worldwide in August 2024 [Patel(2024)], with its audience primarily aged 25-30 [Christopher(2023)].
  • Character.ai: Established in late 2022, this platform allows users to role-play using constructed personas and interact with chatbots based on fictional characters. With over 20 million users reported in March 2024, its audience primarily comprises young adults aged 18-24 [Shewale(2024)]. Notably, the sixth most popular character on the platform is “Psychologist,” with over 70 million messages shared in a year, indicating a significant trend of users seeking mental health support from AI [Tidy(2024)].
  • Snapchat’s My AI: A chatbot integrated into the Snapchat platform, powered by large language models like GPT-3.5 and GPT-4. It offers advice, answers questions, and assists with various tasks, demonstrating the integration of AI companions into mainstream social media platforms. Over 150 million individuals have used My AI since its launch [Hutchinson(2023)].
  • Anima: A chatbot service that enables role-playing conversations with different personas, allowing users to explore various scenarios and characters.
  • Pi: Developed by Inflection AI, Pi is an AI companion designed to engage in open-ended conversations on a wide range of topics. It aims to be a supportive, empathetic presence for users, offering advice, emotional support, and intellectual discussions, with six million monthly active users reported in March 2024 [Wodecki(2024)].
@@
The popularity of these services, particularly among young adults, speaks to their growing acceptance as sources of social interaction and emotional support [Koulouri et al.(2022), Xygkou et al.(2024), Xygkou et al.(2023)]. Research suggests that artificial companions can reduce loneliness through various mechanisms, including direct companionship, acting as a catalyst for social interaction, facilitating remote communication, and providing reminders for social activities [Gasteiger et al.(2021)].
@@
The perceived “realness” of the connection with AI companions is a crucial factor in their effectiveness. Users who perceive chatbots as having human-like qualities, such as consciousness or agency, often experience improved social health and reduced loneliness [Xia et al.(2024)]. This anthropomorphism can enhance the effectiveness of companion chatbots in mitigating feelings of isolation, though it carries a risk of leading to the user dehumanizing other people [Zehnder et al.(2021)]. Pataranutaporn & Liu et al. [Pataranutaporn et al.(2023)] also found that mental models altered experience – priming users with different beliefs about an AI’s motives (caring, manipulative, or neutral) significantly affected their assessment of the AI’s trustworthiness, empathy, and effectiveness. This suggests that users’ mental models of AI systems play a crucial role in shaping their interactions and experiences.
Previous work has suggested that the ability to empathize is inversely correlated with loneliness over a lifetime [Beadle et al.(2012)], but the question remains of what can be the role of a chatbot that simply performs empathy. Role-play has been shown to increase empathy [Bearman et al.(2015), Dewi et al.(2019)], and the potential of AI-facilitated role-play to influence empathy and, by extension, loneliness, represents an unexplored area in the research literature, despite the large number of users engaging in such interactions.
@@
Existing research on the influence of artificial agents on socialization with other people has produced mixed results, with the work mostly qualitative in nature [Van der Loos(2014), Turkle(2011)]. One study suggested that artificial agents may end up replacing humans for task-oriented interactions, while in social contexts they are more likely to complement human social interactions [Xia et al.(2024)].
@@ 2.4 Human-AI Interaction for Emotional Support
Recent studies suggest that even limited amounts of AI companionship reduce the experience of loneliness [De Freitas et al.(2024)]. However, the long-term effects and potential risks of relying on AI for emotional support remain subjects of ongoing investigation and ethical debate [Li et al.(2023), Abd-Alrazaq et al.(2020)].
Skjuve et al. described how human-chatbot relationships (HCRs) evolve from an initial superficial phase through more affective exploration – with greater trust and self-disclosure – to a more stable state of emotional connection [Skjuve et al.(2021)]. This relationship development enables companion chatbots to provide various forms of social support, including companionship, emotional support, and informational guidance in a “safe space” that is always available but not physical [Ta et al.(2020)]. The availability and perceived safety of these interactions make them an attractive alternative for social support for those who may lack support from family and friends, such as LGBTQ individuals [Ma et al.(2024)]. The motivations behind these interactions are diverse, ranging from curiosity to explicit seeking of emotional support, with users engaging in topics from casual conversation to deep personal issues. These findings suggest that while chatbots can serve as valuable sources of emotional support, the nature and impact of these interactions vary significantly based on individual user characteristics and needs [Ta-Johnson et al.(2022)]. Though limited to one service, Siemon et al.’s analysis of Replika reviews [Siemon et al.(2022)] showed that users of Replika often use the service to cope with their loneliness, and their long-term interactions with Replika can be joyful and beneficial.
Recent studies by Jo et al. [Jo et al.(2023), Jo et al.(2024)] explored the deployment of an LLM-driven chatbot, CareCall, for public health interventions targeting socially isolated individuals. Their findings suggest that chatbots can alleviate public health workloads and mitigate loneliness, with users perceiving the system as personal and emotionally supportive. However, challenges such as trust, privacy, and the limitations of LLM-driven chatbots in addressing complex health needs were also highlighted in this study, emphasizing the importance of careful system design and managing stakeholder expectations.
Age can also affect how users interact with AI. Neves et al. [Neves et al.(2015)] examined the difficulties faced by older adults in adopting new communication technologies, highlighting the importance of user-centered design and support systems in facilitating technology acceptance. These insights are crucial when considering the potential of AI chatbots as a scalable solution for emotional support and loneliness reduction.
While companion chatbots offer promising avenues for emotional support, there are significant ethical considerations and potential risks associated with their use. These include the possibility of over-reliance on AI, consequences for mood and affect, delays in seeking professional help, and further withdrawal from human socialization [Pataranutaporn et al.(2021)]. A recent case of a teen suicide after extensive use of Character.ai raised concerns regarding whether or not the chatbot could be blamed for his death [Roose(2024)]; this case underscores the urgent need to understand the influence of companion chatbots on psychosocial well-being. The development of dependence on AI agents and the potential replacement of real human relationships raise important questions about the long-term societal impacts of widespread AI companion use.
More specific, granular information about users and their needs is needed. Our research seeks to identify the characteristics of users who benefit from chatbot interactions and users who may experience negative consequences. Our intent is to inform more targeted and ethical design practices in the rapidly evolving landscape of human-AI interaction towards the goal of positive computing in human-AI interaction – designing technology that supports psychological well-being [Calvo and Peters(2017)].

### v3

@@ Abstract
Companion chatbots offer a potential solution to the growing epidemic of loneliness, but their impact on users’ psychosocial well-being remains poorly understood, raising critical ethical questions about their deployment and design. This study presents a large-scale survey (n = 404) of regular users of companion chatbots, investigating the relationship between chatbot usage and loneliness. We develop a model explaining approximately 50% of variance in loneliness; while usage does not directly predict loneliness, we identify factors including neuroticism, social network size, and problematic use. Through cluster analysis and mixed-methods thematic analysis combining manual coding with automated theme extraction, we identify seven distinct user profiles demonstrating that companion chatbots can either enhance or potentially harm psychological well-being depending on user characteristics. Different usage patterns can lead to markedly different outcomes, with some users experiencing enhanced social confidence while others risk further isolation. These findings have significant implications for responsible AI development, suggesting that one-size-fits-all approaches to AI companionship may be ethically problematic. Our work contributes to the ongoing dialogue about the role of AI in social and emotional support, offering insights for developing more targeted and ethical approaches to AI companionship that complement rather than replace human connections.
Overview of the study’s methodology and major findings. A survey of 404 participants explored chatbot usage and its psychological impacts. From the findings, we developed a multiple regression model and identified seven user archetypes.
@@ 1 Introduction
In an increasingly digital world, we are paradoxically more connected and isolated than ever. Despite unprecedented technological means to connect, loneliness remains pervasive, affecting approximately a third of individuals in industrialized countries [1]. This phenomenon has led to what the U.S. Surgeon General declared an “epidemic of loneliness” [2]. The gravity of this issue is underscored by research indicating that chronic loneliness can increase mortality risk by 26-29%, comparable to smoking 15 cigarettes a day [3].
Artificial intelligence (AI) has emerged as a potential tool for addressing loneliness. Companion chatbots, which engage in complex conversations and maintain persistent relationships with users, have gained significant popularity, with platforms like Replika and Character.ai reporting over 30 million and 20 million users, respectively [4, 5]. These companion chatbots offer users the opportunity for constant, judgment-free interaction. Some research suggests that companion chatbots, even when used over the short term, decrease feelings of loneliness [1, 6, 7]. However, concerns have been raised about the authenticity of these relationships, suggesting that they lower the bar for what people think relationships can be [8]. This could lead to unrealistic expectations and further social withdrawal.
The nuanced nature of human-AI interaction in this context presents a critical gap in our understanding. Specifically, we lack comprehensive knowledge about factors influencing companion chatbots’ impact on loneliness and other psychological and social factors. Artificial agents may complement human social interactions in some contexts while replacing them in others [9]. This variability in outcomes underscores the need for a more comprehensive and granular analysis of user characteristics, usage patterns, and their relationships to psychological well-being. To address this gap, we conducted a survey (n = 404) of regular companion chatbot users, with the following research questions:
@@
Our study aims to provide a more nuanced understanding of the impact of companion chatbots. By identifying the characteristics of users who experience psychosocial benefits from chatbot interactions (i.e., reduced loneliness and improved social skills) and those who may experience negative consequences (i.e., emotional dependence and social withdrawal), we hope to inform more targeted and ethical design practices. For example, more information could guide the development of customizable interfaces that emphasize features beneficial to specific user groups or implement safeguards for at-risk users. By examining how users’ mental models and expectations of companion chatbots influence outcomes, we can explore strategies for fostering healthy human-AI relationships that complement rather than replace human connections.
Through a combination of statistical analyses, including multiple regression and cluster analysis, along with qualitative assessment of user responses, we present the following contributions: (1) A model explaining approximately 50% of variance in loneliness, identifying problematic use as a key mediator and social attraction as a moderator in the relationship between chatbot usage and loneliness while accounting for personality traits and social network characteristics; (2) a typology of seven distinct companion chatbot user profiles demonstrating how similar usage patterns can lead to markedly different outcomes based on individual characteristics and social contexts; (3) an exploration of motivations and usage characteristics across different user types, revealing how technological exploration and entertainment dominate over social purposes despite marketing as social tools; (4) a demonstration of how manual thematic analysis can be effectively triangulated with LLM-based theme extraction to achieve both interpretive depth and comprehensive coverage in large-scale qualitative analysis; and (5) design implications for creating more effective and ethically responsible companion chatbots, including the need for personalized approaches and built-in mechanisms to detect and intervene in potentially harmful usage patterns.
@@ 2.1 Loneliness and Social Support
Loneliness, the subjective feeling of isolation regardless of actual social network, affects approximately one-third of individuals in industrialized countries [1]. The U.S. Surgeon General’s 2023 declaration of loneliness as an epidemic highlights this issue’s significance [2]. Chronic loneliness increases mortality risk by 26-29%, comparable to smoking 15 cigarettes daily [10].
Individuals employ various coping strategies for loneliness [11, 12], with digital platforms emerging as common mechanisms. Evidence shows that lonely, socially isolated, and socially anxious individuals tend to engage in one-sided paradigms like parasocial relationships with content creators [13, 14, 15, 16] and face a higher risk of addiction to online platforms [17, 18, 19].
Effective interventions for reducing loneliness include cognitive behavioral therapy, improving social support, and facilitating social interaction [11, 6, 20], though efficacy varies across individuals. This aligns with recognition that lonely individuals constitute a diverse group with varying needs [21], suggesting interventions—including technological ones—must be tailored to individual circumstances.
@@ 2.2 Companion Chatbots
Companion chatbots represent a novel approach to addressing loneliness [22, 23, 24, 25, 26, 27]. These commercially available chatbots employ generative AI to engage in complex conversations and maintain persistent relationships with users. Several popular services have gained significant user bases: Replika reported over 30 million registered users worldwide by August 2024 [4], while Character.ai, a platform allowing users to role-play with chatbots based on fictional characters, reported over 20 million users in March 2024 [5].
@@
The popularity of these services, particularly among young adults, indicates their growing acceptance as sources of social interaction and emotional support [28, 29, 30]. Research suggests artificial companions can reduce loneliness through direct companionship, acting as catalysts for social interaction, facilitating remote communication, and providing reminders for social activities [6].
@@
The perceived “realness” of connection with companion chatbots significantly impacts their effectiveness. Users who attribute human-like qualities to chatbots often experience improved social health and reduced loneliness [9], though this carries a risk of dehumanizing other people [31]. Pataranutaporn & Liu et al. [32] found that mental models altered experience—priming users with different beliefs about an AI’s motives significantly affected their assessment of the AI’s trustworthiness and effectiveness.
@@
Research on artificial agents’ influence on human socialization has produced mixed results, mostly qualitative in nature [33, 8]. One study suggested artificial agents may replace humans for task-oriented interactions while complementing human social interactions [9].
While empathy is inversely correlated with loneliness [34], the role of chatbots that perform empathy remains uncertain. Role-play increases empathy [35, 36], yet the potential of AI-facilitated role-play to influence empathy and loneliness represents an unexplored research area despite widespread user engagement.
@@ 2.3 Human-AI Interaction for Emotional Support
Studies suggest even limited AI companionship reduces loneliness [7], though long-term effects and risks remain subjects of investigation and ethical debate [37, 38]. Skjuve et al. described how human-chatbot relationships evolve from superficial interaction through affective exploration to stable emotional connection [39]. This enables companion chatbots to provide various forms of social support in an always-available “safe space” [40]. These interactions appeal particularly to those lacking traditional support systems, such as LGBTQ individuals [41]. User motivations range from curiosity to explicit seeking of emotional support, with topics spanning casual conversation to deep personal issues [42]. Siemon et al.’s analysis of Replika reviews [43] showed users often employ the service to cope with loneliness, with long-term interactions proving beneficial.
Significant ethical considerations include potential over-reliance on AI, consequences for mood, delays in seeking professional help, and withdrawal from human socialization [44]. A teen suicide case in 2024 following extensive Character.ai use raised concerns about chatbot influence on mental health [45], underscoring the need to understand companion chatbots’ impact on psychosocial well-being.
Our research seeks to identify characteristics of users who benefit from chatbot interactions versus those who experience negative consequences, informing targeted and ethical design practices toward positive computing in human-AI interaction—designing technology that supports psychological well-being [46].